does anybody know how to solve y=x^2+x-6
Answer by ewatrrr(24785)


Hi
Solving for x when y = 0
x^2 + x-6 = 0
factoring
(x+3)(x-2)= 0 Note:SUM of the inner product(3x) and the outer product(-2x) = x
(x+3)=0
x = -3
(x-2)= 0
x = 2